An object starts with a velocity that has an x-component of -9m/s and y-component of 8m/s. It ends with a velocity that has an x-component of 7m/s and y-component of -4m/s. What are the components its displacement after 6.5sec of moving with a constant acceleration?
Answer #1

From kinematic equation

Vx =Vxo+ax*t

7 m/s =-9 m/s +ax*6.5 s

ax=2.46 m/s^2

Vx^2 =Vxo^2+2*ax*X

(7 m/s)^2 =(-9 m/s)^2 +2*2.46 m/s^2*X

X= -6.50 m

Vy=Vyo +ay*t

ay =-1.846 m/s^2

Vy^2=Vyo^2+2*ay*Y

Y = 13 m

the components of the displacement is (X,Y)=(-6.50 ,13) m